After robbing a bank in Dodge City, a robber gallops off at 12 mi/h. 10 minutes later, the marshall leaves to pursue the robber at 16 mi/h. How long (in hours) does it take the marshall to catch up to the robber?

Answer by stanbon(75887)

After robbing a bank in Dodge City, a robber gallops off at 12 mi/h. 10 minutes later, the marshall leaves to pursue the robber at 16 mi/h. How long (in hours) does it take the marshall to catch up to the robber?
-------------------
Robber DATA:
rate = 12 mph ; time = x hrs ; distance = 12x miles
--------------
Marshall DATA:
rate = 16 mph ; time = x-(1/6) hr distance = 16(x-(1/6)) miles
----
Equation:
distance = distance
12x = 16(x-(1/6))
12x = 16x-(16/6)
-4x = -(16/6)
x = 2/3 hr
----
x-(1/6) = (4/6)-(1/6) = 1/2 hr (time needed for Marshall to catch robber)
==============
Cheers,
Stan H.